# Computation method for perfectoid purity and perfectoid BCM-regularity

### v4

@@ 4.1. Definition of quasi-$F$-splitting
We use Notation 4.2. We assume $R$ is $p$-torsion free. For positive integers $n,m$, we obtain the following exact sequence as $\overline{R}$-modules:
@@
First, let us show that $V^{m}\colon F^{m}_{*}Q_{R,n}\to Q_{R,n+m}$ is an $\overline{R}$-module homomorphism. We take $\alpha\in W_{n}(R)$ and $b\in R$. Then we have $[b^{p}]V^{m}\alpha=V([b^{p^{m+1}}]\alpha)$ by Proposition 2.3, thus $V^{m}$ is an $\overline{R}$-module homomorphism. Next, let us show the exactness. It is enough to show that the sequence
@@
(cf. [Yobuko19]) We use Notation 4.2. We assume $R$ is Noether and $\overline{R}$ is $F$-finite.
@@
We use Notation 4.2. We assume $R$ is Noetherian $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\mathrm{ht}(R)\leq\mathrm{ht}(\overline{R})$. Furthermore, if $R$ is normal and $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})=\operatorname{ht}^{e}(\overline{R})$ for every $e\geq 1$, where $\operatorname{ht}^{e}(\overline{R})$ is defined in [TWY]*Definition 3.5, then $R$ is quasi-$(F,F^{\infty})$-split.
@@
We use Notation 4.2 and assume $(R,\mathfrak{m})$ is Noether local $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\operatorname{ht}(R)=\operatorname{ht}(\widehat{R})$, where $\widehat{R}$ is the $m$-adic completion of $R$. Furthermore, $R$ is quasi-$(F,F^{\infty})$-split if and only if so is $\widehat{R}$.
@@
We use Notation 4.2 and assume $R$ is $p$-torsion free. If $R$ is Cohen-Macaulay, then so is $Q_{R,n}$ for every positive integer $n$.
@@
Let us show the assertion by induction on $n$. Since $R$ is Cohen-Macaulay and $p$-torsion free, the ring $\overline{R}\simeq W_{R,1}$ is Cohen-Macaulay. We take an integer $n\geq 2$. By Proposition 4.3, we obtain an exact sequence
@@
We use Notation 4.2 and assume $(R,\mathfrak{m})$ is local. Let $f\in\mathfrak{m}$ be a regular element such that $R/f$ is $p$-torsion free. We assume $R$ is Gorenstein. Then $(R,f)$ is purely quasi-$F$-split if and only if $R/f$ is quasi-$F$-split.
@@ 4.2. Quasi-$F$-splitting for perfectoids
In this subsection, we study the purity of $\Phi_{R,n}$ for a perfectoid $R$. Theorem 4.10 is related to the computation of perfectoid pure threshold (cf. Theorem 5.8).
@@
Let $R^{\flat}$ be a tilt of $R$ and $T\in R^{\flat}$ the element corresponding to a system $\{p^{1/p^{e}}\}$. Let $A:=W(R^{\flat})$ and $d:=[T]-p$, then $(A,(d))$ is a perfect prism with $A/(d)\simeq R$ by [BS]*Corollary 2.31. The lift of Frobenius on $A$ is denoted by $\phi$. Let us show the following claim:
@@
Since $\delta_{1}(d)$ is a unit, the map $\psi_{1}$ satisfies the condition (a). On the other hand, we have
@@
thus $\psi_{1}$ also satisfies the condition (b). Furthermore, we have $\psi(a)=F^{-1}(a)$, thus $\psi$ satisfies the condition (c). Next, we assume $n\geq 2$. By the induction hypothesis, there exists $\psi_{n-1}\colon Q_{A,n-1}\to R^{\flat}$ such that $\psi_{n-1}$ satisfies the conditions (a), (b), (c). We replace $\psi_{n-1}$ by $(-\psi_{n-1}(\Delta_{W_{n}}(d))^{-1})\cdot\psi_{n-1}$, we may assume that $\psi_{n-1}(\Delta_{W_{n}}(d))=-1$. We set
@@
Thus, the homomorphism $\psi_{n}$ satisfies the condition (b). Furthermore, we have $\psi_{n}(1)=F^{-1}(1)=1$ and
@@
thus, the map $\psi_{n}$ satisfies the conditions (a), (b), (c). $\blacksquare$
We take $\psi_{n}$ as in Claim 4.11. Replacing $\psi_{n}$ by $(\psi_{n}(1))^{-1}\cdot\psi_{n}$, we may assume $\psi_{n}(1)=1$. By the condition (b), $\psi_{n}$ induces the map
@@
with $\overline{\psi}_{n}(1)=1$. Furthermore, by the condition (c), there exists a unit $w\in\overline{R}$ such that $\overline{\psi}(V^{n-1}[a])=wT^{i_{n}}F^{-n}(a)$ for every $a\in A$. ∎
4.3. Fedder type criterion for quasi-$F$-splitting
In this subsection, we introduce Fedder type criterion (Theorem 4.12) for quasi-$F$-splitting and quasi-$(F,F^{\infty})$-splitting. The proof of Theorem 4.12 is almost identical to a proof of [TWY]*Theorem A, however, for the convenience of the reader, we provide a proof in Appendix C. Furthermore, we introduce examples of computations of quasi-$F$-splitting height and quasi-$(F,F^{\infty})$-splitting by using Theorem 4.12.
@@ 4.3. Fedder type criterion for quasi-$F$-splitting
    We set $I^{\prime}:=\bigcap_{e\geq 1}u^{e}(F^{e}_{*}(f^{p^{e}-1}\overline{A}))$. We define a sequence $\{I^{\prime}_{n}\}$ of ideals of $\overline{A}$ by $I^{\prime}_{1}:=f^{p-1}I^{\prime}+I^{[p]}$ and
    $I^{\prime}_{n}:=u(F_{*}(\Delta_{1}(f^{p-1})I^{\prime}_{n-1}))+f^{p-1}\overline{A}+I^{[p]}$
@@
    Let $p=2,N=3,f=x^{3}+y^{3}+z^{3}$. Then $I_{1}\subseteq\mathfrak{m}^{[p]}$ and $u(F_{*}(f\Delta_{1}(f))\equiv xyz\ \textrm{mod}\ \mathfrak{m}^{[p]}$, thus we have $I_{2}\nsubseteq\mathfrak{m}^{[p]}$. Furthermore, we have $u(F_{*}(f\Delta_{1}(f)\mathfrak{m}))\subseteq\mathfrak{m}^{[p]}$, thus we have $I_{2}^{\prime}\subseteq\mathfrak{m}^{[p]}$, in conclusion, $R$ is not quasi-$(F,F^{\infty})$-split (cf. Proposition 6.8).
@@
    Let $p=2,N=3,f=z^{2}+x^{2}y+xy^{n}$ for $n\geq 2$. First, we note that $z\in I^{\prime}$. By the argument in [kty]*Section 7.1, we have $\theta^{m}(F^{m}_{*}(zy^{2^{m}-n}f))=xyz\notin\mathfrak{m}^{[p]}$ and $\operatorname{ht}(R)=m+1$, where $m:=\lceil\log_{2}n\rceil$, thus $R$ is quasi-$(F,F^{\infty})$-split.
@@
$\Delta_{1}((ff^{\prime})^{p-1})=f^{p(p-1)}\Delta_{1}(f^{\prime p-1})+f^{\prime p(p-1)}\Delta_{1}(f^{p-1})\in\mathfrak{m}^{[p^{2}]},$
@@
5. Quasi-$F$-splitting versus perfectoid purity
In this section, we study the relationship between quasi-$F$-splitting and perfectoid purity. For this purpose, we introduce a program (Definition 5.4) to obtain a sequence of elements of local cohomologies. Whether or not this program stops is closely related to the singularity of a given ring.
@@ 5. Quasi-$F$-splitting versus perfectoid purity
We use Notation 5.1. Let $n$ be a positive integer and $i_{n}=1/p+\cdots+1/p^{n}$. We assume $R$ is Gorenstein. Let $\rho\colon R\to R_{\infty}$ be a ring homomorphism to a $p$-torsion free perfectoid $R_{\infty}$. If $\operatorname{ht}(R)\geq n+1$, then
@@
We use Notation 5.1. We assume $R$ is Gorenstein. If $R$ is not quasi-$F$-split and $p=2$, then $R$ is not perfectoid BCM-regular.
@@
We use Notation 5.1. We take $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and consider the following program. We set $\eta_{0}:=\eta$ and assume $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ is defined for $i\geq 0$.
  1. Case(i):
    We consider the case of $F(\eta_{i})\neq 0$, where $F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(F_{*}\overline{R})$.
    1. Case(i-i):
      If there exists $e^{\prime}\geq 2$ such that $F^{e^{\prime}}(\eta_{i})=0$, then we take the integer $e\geq 1$ such that $F^{e}(\eta_{i})\neq 0$ and $F^{e+1}(\eta_{i})=0$ and we set $\eta_{i+1}:=F^{e}(\eta_{i})$.
    2. Case(i-ii):
      If $F^{e}(\eta_{i})\neq 0$ for every integer $e\geq 2$, then we stop the program.
  2. Case(ii):
    We consider the case of $F(\eta_{i})=0$.
    1. Case(ii-i):
      If there exists a positive integer $n^{\prime}\geq 2$ such that $\Phi_{R,n^{\prime}}(\eta)\neq 0$, where $\Phi_{R,n^{\prime}}\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(Q_{R,n^{\prime}})$, then we take an integer $n\geq 2$ such that $\Phi_{R,n-1}(\eta)=0$ and $\Phi_{R,n}(\eta)\neq 0$. By the exact sequence in Proposition 4.3, there exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})$ such that $V^{n-1}(\eta^{\prime})=\Phi_{R,n}(\eta_{i})$. Then we chose such $\eta^{\prime}$ and we set $\eta_{i+1}:=\eta^{\prime}$.
    2. Case(ii-ii):
      If $\Phi_{R,n}(\eta_{i})=0$ for every $n\geq 2$, then we stop the program.
Remark 5.5.
Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$.
  1. (1)
    In the program in Definition 5.4, if $\eta_{i}$ is in the case (i-i), then $\eta_{i+1}$ is in the case (ii).
  2. (2)
    The program in Definition 5.4 has the following three outputs.
    • We can define $\eta_{0},\eta_{1},\ldots,\eta_{i}$ for some $i\geq 0$ and $\eta_{i}$ is in the case (i-ii).
    • We can define $\eta_{0},\eta_{1},\ldots,\eta_{i}$ for some $i\geq 0$ and $\eta_{i}$ is in the case (ii-ii).
    • We can define $\eta_{i}$ for all $i\geq 0$.
  3. (3)
    If $R$ is quasi-$F$-split, the case (ii-ii) does not occur since $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(Q_{R,n})$ is injective for some $n\geq 1$.
@@ 2. Case(ii-ii):
We use Notation 5.1. Let $S$ be an $R$-algebra and $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$. We assume $S$ is $p$-torsion free and $H^{d-1}_{\mathfrak{m}}(\overline{S})=0$, where $\overline{S}:=S/pS$. We run a program in Definition 5.4 for $\eta_{0}:=\eta$ and get $\eta_{i}$ in the $i$-th step. If $\eta$ is contained in $\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{S}))$, then so is $\eta_{i}$.
@@
It is enough to show $\eta_{1}\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{S}))$. If $\eta$ is in the case (i-i), then $\eta_{1}=F^{e}(\eta)$ for some $e\geq 1$ and the assertion follows from the commutative diagram

### v5

@@ 4.1. Definition of quasi-$F$-splitting
We use the notation introduced in Notation 4.2. We assume $R$ is $p$-torsion free. For positive integers $n,m$, we obtain the following exact sequence as $\overline{R}$-modules:
@@
First, we prove that $V^{m}\colon F^{m}_{*}Q_{R,n}\to Q_{R,n+m}$ is an $\overline{R}$-module homomorphism. We take $\alpha\in W_{n}(R)$ and $b\in R$. Then we have $[b^{p}]V^{m}\alpha=V([b^{p^{m+1}}]\alpha)$ by Proposition 2.3, thus $V^{m}$ is an $\overline{R}$-module homomorphism. Next, we prove the exactness. It is enough to show that the sequence
@@
(cf. [Yobuko19]) We use the notation introduced in Notation 4.2. We assume $R$ is Noether and $\overline{R}$ is $F$-finite.
@@
We use the notation introduced in Notation 4.2. We assume $R$ is Noetherian $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\mathrm{ht}(R)\leq\mathrm{ht}(\overline{R})$. Furthermore, if $R$ is normal and $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})=\operatorname{ht}^{e}(\overline{R})$ for every $e\geq 1$, where $\operatorname{ht}^{e}(\overline{R})$ is defined in [TWY]*Definition 3.5, then $R$ is quasi-$(F,F^{\infty})$-split.
@@
We use the notation introduced in Notation 4.2 and assume $(R,\mathfrak{m})$ is Noether local $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\operatorname{ht}(R)=\operatorname{ht}(\widehat{R})$, where $\widehat{R}$ is the $m$-adic completion of $R$. Furthermore, $R$ is quasi-$(F,F^{\infty})$-split if and only if so is $\widehat{R}$.
@@
We use the notation introduced in Notation 4.2 and assume $R$ is $p$-torsion free. If $R$ is Cohen-Macaulay, then so is $Q_{R,n}$ for every positive integer $n$.
@@
We prove the assertion by induction on $n$. Since $R$ is Cohen-Macaulay and $p$-torsion free, the ring $\overline{R}\simeq W_{R,1}$ is Cohen-Macaulay. We take an integer $n\geq 2$. By Proposition 4.3, we obtain an exact sequence
@@
We use the notation introduced in Notation 4.2 and assume $(R,\mathfrak{m})$ is local. Let $f\in\mathfrak{m}$ be a regular element such that $R/f$ is $p$-torsion free. We assume $R$ is Gorenstein. Then $(R,f)$ is purely quasi-$F$-split if and only if $R/f$ is quasi-$F$-split.
@@ 4.2. Quasi-$F$-splitting for perfectoids
In this subsection, we study the purity of $\Phi_{R,n}$ for a perfectoid $R$. Theorem 4.10 is related to the computation of perfectoid pure threshold (cf. LABEL:ht-to-p-pure).
@@
Let $R^{\flat}$ be a tilt of $R$ and $T\in R^{\flat}$ the element corresponding to a system $\{p^{1/p^{e}}\}$. Let $A:=W(R^{\flat})$ and $d:=[T]-p$, then $(A,(d))$ is a perfect prism with $A/(d)\simeq R$ by [BS]*Corollary 2.31. The lift of Frobenius on $A$ is denoted by $\phi$. We prove the following claim:
@@
Since $\delta_{1}(d)$ is a unit, the map $\psi_{1}$ satisfies condition (a). On the other hand, we have
@@
thus $\psi_{1}$ also satisfies condition (b). Furthermore, we have $\psi(a)=F^{-1}(a)$, thus $\psi$ satisfies condition (c). Next, we assume $n\geq 2$. By the induction hypothesis, there exists $\psi_{n-1}\colon Q_{A,n-1}\to R^{\flat}$ such that $\psi_{n-1}$ satisfies conditions (a), (b), (c). We replace $\psi_{n-1}$ by $(-\psi_{n-1}(\Delta_{W_{n}}(d))^{-1})\cdot\psi_{n-1}$, we may assume that $\psi_{n-1}(\Delta_{W_{n}}(d))=-1$. We set
@@
Thus, the homomorphism $\psi_{n}$ satisfies condition (b). Furthermore, we have $\psi_{n}(1)=F^{-1}(1)=1$ and
@@
thus, the map $\psi_{n}$ satisfies conditions (a), (b), (c). $\blacksquare$
We take $\psi_{n}$ as in Claim 4.11. Replacing $\psi_{n}$ by $(\psi_{n}(1))^{-1}\cdot\psi_{n}$, we may assume $\psi_{n}(1)=1$. By condition (b), $\psi_{n}$ induces the map
@@
with $\overline{\psi}_{n}(1)=1$. Furthermore, by condition (c), there exists a unit $w\in\overline{R}$ such that $\overline{\psi}(V^{n-1}[a])=wT^{i_{n}}F^{-n}(a)$ for every $a\in A$. ∎
4.3. Fedder-type criterion for quasi-$F$-splitting
In this subsection, we introduce Fedder-type criterion (Theorem 4.12) for quasi-$F$-splitting and quasi-$(F,F^{\infty})$-splitting. The proof of Theorem 4.12 is almost identical to a proof of [TWY]*Theorem A, however, for the convenience of the reader, we provide a proof in Appendix C. Furthermore, we introduce examples of computations of quasi-$F$-splitting height and quasi-$(F,F^{\infty})$-splitting by using Theorem 4.12.
@@ 4.3. Fedder-type criterion for quasi-$F$-splitting
    We set $I^{\prime}:=\bigcap_{e\geq 1}u^{e}\left(F^{e}_{*}(f^{p^{e}-1}\overline{A})\right)$. We define a sequence $\{I^{\prime}_{n}\}$ of ideals of $\overline{A}$ by $I^{\prime}_{1}:=f^{p-1}I^{\prime}+I^{[p]}$ and
    $I^{\prime}_{n}:=u\left(F_{*}(\Delta_{1}(f^{p-1})I^{\prime}_{n-1})\right)+f^{p-1}\overline{A}+I^{[p]}$
@@
    Let $p=2,N=3,f=x^{3}+y^{3}+z^{3}$. Then $I_{1}\subseteq\mathfrak{m}^{[p]}$ and $u\left(F_{*}(f\Delta_{1}(f)\right)\equiv xyz\ \textrm{mod}\ \mathfrak{m}^{[p]}$, thus we have $I_{2}\nsubseteq\mathfrak{m}^{[p]}$. Furthermore, we have $u\left(F_{*}(f\Delta_{1}(f)\mathfrak{m})\right)\subseteq\mathfrak{m}^{[p]}$, thus we have $I_{2}^{\prime}\subseteq\mathfrak{m}^{[p]}$, in conclusion, $R$ is not quasi-$(F,F^{\infty})$-split (cf. Proposition 6.8).
@@
    Let $p=2,N=3,f=z^{2}+x^{2}y+xy^{n}$ for $n\geq 2$. First, we note that $z\in I^{\prime}$. By the argument in [kty]*Section 7.1, we have $\theta^{m}\left(F^{m}_{*}(zy^{2^{m}-n}f)\right)=xyz\notin\mathfrak{m}^{[p]}$ and $\operatorname{ht}(R)=m+1$, where $m:=\lceil\log_{2}n\rceil$, thus $R$ is quasi-$(F,F^{\infty})$-split.
@@
$\Delta_{1}\left((ff^{\prime})^{p-1}\right)=f^{p(p-1)}\Delta_{1}(f^{\prime p-1})+f^{\prime p(p-1)}\Delta_{1}(f^{p-1})\in\mathfrak{m}^{[p^{2}]},$
@@
5. Quasi-$F$-splitting versus perfectoid purity and perfectoid pure threshold
In this section, we study the relationship between quasi-$F$-splitting and perfectoid purity. For this purpose, we introduce a procedure (Definition 5.4) to obtain a sequence of elements of local cohomologies.
@@ 5.1. Quasi-$F$-splitting versus perfectoid purity
We use the notation introduced in Notation 5.1. Let $n$ be a positive integer and $i_{n}=1/p+\cdots+1/p^{n}$. We assume $R$ is Gorenstein. Let $\rho\colon R\to R_{\infty}$ be a ring homomorphism to a $p$-torsion free perfectoid $R_{\infty}$. If $\operatorname{ht}(R)\geq n+1$, then
@@
We use the notation introduced in Notation 5.1. We assume $R$ is Gorenstein. If $R$ is not quasi-$F$-split and $p=2$, then $R$ is not perfectoid BCM-regular.
@@
We use the notation introduced in Notation 5.1 and assume that $R$ is quasi-$F$-split. We take $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and construct sequences $\{\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}\}_{i\geq 0}$ and $\{n_{i}\in\mathbb{Z}_{\geq 1}\}_{i\geq 0}$ using the following procedure. We start with $\eta_{0}:=\eta$. Suppose that $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ has been defined for some $i\geq 0$. We take the minimal integer $n_{i}\geq 1$ such that $\Phi_{R,n_{i}}(\eta_{i})\neq 0$. We note that such an integer $n_{i}$ exists since $R$ is quasi-$F$-split and $n_{i}\leq\operatorname{ht}(R)$. By the exact sequence in Proposition 4.3, there exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})$ such that $V^{n_{i}-1}(\eta^{\prime})=\Phi_{R,n_{i}}(\eta_{i})$. We choose such an $\eta^{\prime}$ and set $\eta_{i+1}:=\eta^{\prime}$.
Theorem 5.5.
Proof.
We set $\overline{A}:=A/(p)$, $\overline{R}:=R/(p)$, $\overline{R}_{\infty}:=R_{\infty}/(p)$, $\overline{A}_{\infty}:=A_{\infty}/(p)$, and $d:=\dim\overline{R}$. We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. We construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ by the procedure in Definition 5.4. We fix a system of $p$-power roots $\{f^{1/p^{e}}\}$ of $f$ in $A_{\infty}$. By the exact sequence
and $\tau_{i}^{(e)}:=\tau_{i}f^{1-1/p^{e}}$ for all integers $e,i\geq 0$.
Claim 5.6.
We have $\tau_{i}^{(e)}\neq 0$ for all integers $e,i\geq 0$.
Proof.
We prove the assertion by induction on $e$. Since $A$ is Cohen-Macaulay, the homomorphism $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(\overline{A})$ is injective. Since $A\to A_{\infty}$ is pure, the homomorphism $H^{d+1}_{\mathfrak{m}}(\overline{A})\to H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty})$ is injective, and in particular, we have $\tau_{i}^{(0)}=\tau_{i}\neq 0$ for every integer $i\geq 0$. Thus, we assume $e\geq 1$ and take $i\geq 0$. First, we assume $n_{i}=1$, then we have $F(\eta_{i})=\eta_{i+1}$. By the commutative diagram
${H^{d}_{\mathfrak{m}}(\overline{R})}$${H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty})}$${H^{d}_{\mathfrak{m}}(F_{*}\overline{R})}$${H^{d+1}_{\mathfrak{m}}(F_{*}\overline{A}_{\infty}),}$$\scriptstyle{u}$$\scriptstyle{F}$$\scriptstyle{\cdot f^{p-1}F}$$\scriptstyle{F_{*}u}$
we have
$\tau_{i+1}^{(e-1)}$ $=u(\eta_{i+1})f^{1-1/p^{e-1}}=u\circ F(\eta_{i})f^{1-1/p^{e-1}}$
$=F\left(u(\eta_{i})\right)f^{p-1}f^{1-1/p^{e-1}}=F\left(u(\eta_{i})f^{1-1/p^{e}}\right)$
$=F(\tau_{i}^{(e)}).$
By the induction hypothesis, we have $\tau_{i+1}^{(e-1)}\neq 0$, thus we have $\tau_{i}^{(e)}\neq 0$, as desired. Next, we assume $n_{i}\geq 2$, then $\Phi_{R,n_{i}}(\eta_{i})=V^{n_{i}-1}(F^{n_{i}}_{*}\eta_{i+1})$. We set $n:=n_{i}$ and take regular sequence $x_{1},\ldots,x_{d}$ on $\overline{R}$ and elements $a,b\in\overline{R}$ such that $\eta_{i}=[a/x_{1}\cdots x_{d}]$, $\eta_{i+1}=[b/x_{1}^{p^{n}}\cdots x_{d}^{p^{n}}]$, and
$[a^{p}]\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{R,n}.$
In particular, we have
By Claim 5.6, we have $\tau f^{1-p^{1/p^{e}}}=\tau_{0}^{(e)}\neq 0$ for every integer $e\geq 0$. Next, we prove that $R\to R_{\infty}$ is pure. By the argument of [p-pure]*Proposition 6.5, it is enough to show that $fA\to(f^{1/p^{\infty}})$ is pure. Since $(f^{1/p^{\infty}})\simeq\mathrm{colim}_{e}f^{1/p^{e}}A_{\infty}$, it is enough to show that $fA\to f^{1/p^{e}}A_{\infty}$ is pure for every $e\geq 0$. We fix an integer $e\geq 0$. Since $f$ is a regular element in $A$ and $A_{\infty}$, it is enough to show that the homomorphism $A\to A_{\infty}\xrightarrow{\cdot f^{1-1/p^{e}}}A_{\infty}$ is pure. We consider the following commutative diagram in which horizontal sequence is exact:
[…]
Theorem 5.7.
Proof.
By Proposition 4.7 and [p-pure]*Lemma 4.13, we may assume that $R$ is complete. Therefore, there exist a regular local ring $R_{0}$ and regular sequence $f_{1},\ldots,f_{r}$ such that $R_{0}/(f_{1},\ldots,f_{r})\simeq R$. Let $R_{i}:=R_{0}/(f_{1},\ldots,f_{i})$ for $r\geq i\geq 1$. We take a perfectoid $R_{0,perfd}$ as in [p-pure]*Section 4.3. By Andre’s flatness lemma [BS]*Theorem 7.14, there exists a $p$-completely faithfully flat extension $R_{0,perfd}\to R_{0,\infty}$ to a perfectoid such that $R_{0,\infty}$ has systems of $p$-power roots of $f_{1},\ldots,f_{r}$. We set $R_{i,\infty}:=(R_{0,\infty}/(f_{1},\ldots,f_{i}))_{perfd}$. We prove that $R_{i}\to R_{i,\infty}$ is pure for each $r\geq i\geq 0$ by induction on $i$. Since $R_{0}\to R_{0,\infty}$ is faithfully flat, we obtain the case of $i=0$. We take $i\geq 1$ and assume $R_{i-1}\to R_{i-1,\infty}$ is pure. By Theorem 5.5, it is enough to show that $R_{i-1,\infty}$ is cohomologically Cohen-Macaulay and $f_{i},p$ are regular elements of $R_{i-1,\infty}$. The cohomologically Cohen-Macaulayness follows from [p-pure]*Theorem 4.24. Furthermore, the sequence $p,f$ is a regular sequence on $R_{i-1,\infty}$ by the proof of [p-pure]*Theorem 4.24, and in particular, the element $p$ is a regular element. By Lemma 5.8, the element $f$ is a regular element of $R_{i-1,\infty}$, as desired. The last assertion follows from the first assertion and Proposition 4.6. ∎
@@
5.2. Computation of perfectoid threshold via quasi-$F$-splitting
Theorem 5.9.
We use the notation introduced in Notation 5.1 and assume that $R$ is quasi-$F$-split and complete intersection. We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. We construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ and $\{n_{i}\}_{i\geq 0}$ using the procedure in Definition 5.4. Then we have
$\operatorname{ppt}(R;{\rm div}(p))=\sum_{m\geq 1}\frac{a_{m}}{p^{m}},$
where
$a_{m}:=\begin{cases}p-1&\text{if $m=\sum_{i=0}^{r}n_{i}$ for some $r\geq 0$}\\ p-2&\text{otherwise}\end{cases}$
Proof.
We take a perfectoid $R_{\infty}$ over $R$ as in [BMPSTWW24]*Lemma 4.23 and set
where
$\eta^{\prime}_{i}:=(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))(\eta_{i})$
for every $i\geq 0$. Then $\operatorname{ppt}(R;{\rm div}(p))=\alpha_{0}$ by Proposition 2.9.
Claim 5.10.
We have
$\alpha_{j}=1-i_{n_{j}}+p^{-n_{j}}\alpha_{j+1}$
for every $j\geq 0$, where
$i_{m}:=\frac{1}{p}+\cdots+\frac{1}{p^{m}}$
for every integer $m\geq 1$.
Proof.
We fix $j\geq 0$ and set $n:=n_{j}$. We take a system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$ in $R_{\infty}$ and $T\in R_{\infty}^{\flat}$ is the corresponding element of the projective system $\{p^{1/p^{e}}\}_{e\geq 0}$. We consider the following commutative diagram;